how much water should be added to 1 gallon of pure antifreeze to obtain a solution of 95% antifreeze?
Found 2 solutions by nerdybill, RAY100:
Answer by nerdybill(7384)

how much water should be added to 1 gallon of pure antifreeze to obtain a solution of 95% antifreeze?
.
Let x = gallons of water added
then
1 = .95(1+x)
1 = .95+.95x
.05 = .95x
.05/.95 = x
.0526 gallons = x (amount of water to add)

Answer by RAY100(1637)

Let x = amt of pure water added (0% antifreeze)
.
(0 %) x + 1(100%) = (1+x)(95%)
.
0*x + 1+1.00 = .95 +.95x
.
1=.95 +.95x
.
.05 = .95x
.
.053 = x,,,, in gallons
.
.
Check
.
0(.053) +1 = .95(1.053)= 1,,,,,ok
.